You have a radioactive sample which has an activity of 3.89×105 Bq at t = 0 s. Later on at t = 1540 s you measure the activity again and find it is 3.15×105 Bq.

What is the half-life of the sample?
1
Expert's answer
2020-05-28T13:10:07-0400

Given, initial activity i.e activity of the radio active element at "t=0s" is "A=A_0=3.89\times 10^5Bq" , and at time "t=1540s" ,activity "A=3.15\times10^5Bq" .


Since, Activity of radio active elements is governed by at any time "t" is given by,


"A=A_0e^{-\lambda t}"

where,"\lambda" is decay constant.


Thus, we get,


"3.15\times10^5=3.89\times10^5e^{-1540\lambda}\\
\implies\frac{3.15}{3.89}=e^{-1540\lambda}\\
\implies \lambda=\frac{1}{1540}ln\bigg(\frac{3.89}{3.15}\bigg)\\
\implies \lambda=1.37\times 10^{-4}s^{-1}"

Now, we also know that, half life of radio active element is given by,


"t_{\frac{1}{2}}=\frac{ln2}{\lambda}\\
\implies t_{\frac{1}{2}}=\frac{0.963}{1.37\times10^{-4}}=5059.4\approx5059s"

Thus,

"\boxed{t_{\frac{1}{2}}=5059s}"